OpenStudy (anonymous):

A delivery service charges a base price for an overnight delivery of a package plus an extra charge for each pound the package weighs. A customer is billed $22.85 for shipping a 3-pound package and $40 for shipping a 10-pound package.

OpenStudy (anonymous):

Write an equation that gives the total cost of shipping a package as a function of the weight of the package.

OpenStudy (ash2326):

Let C be the total charge B= base price x= weight in punds charge per pound =P C=B+Px for 3 pounds C=22.85 a) 22.85=B+3P for 10 pounds C=40 b) 40=B+10P subtracting a from b 17.15=7P P=2.45 per pound 40=B+24.5 B=15.5 so total Charge =15.5+2.45x x=weight in pounds

OpenStudy (anonymous):

Find the cost of shipping a 15-pound package.

OpenStudy (ash2326):

put x=15 pound in charge equation charge=15.5+2.45*15= you find now

OpenStudy (anonymous):

$52.25